# A Simple Approach to Constraint-Aware Imitation Learning with Application to Autonomous Racing

### v1

@@ Abstract
Guaranteeing constraint satisfaction is challenging in imitation learning (IL), particularly in tasks that require operating near a system’s handling limits. Traditional IL methods often struggle to enforce constraints, leading to suboptimal performance in high-precision tasks. In this paper, we present a simple approach to incorporating safety into the IL objective. Through simulations, we empirically validate our approach on an autonomous racing task with both full-state and image feedback, demonstrating improved constraint satisfaction and greater consistency in task performance compared to a baseline method.
@@ II-A System
As a shorthand notation, we denote $\mathbf{u}_{0:N-1}=\{u_{0},\dots,u_{N-1}\}$ as a control sequence. The state at time $N$ resulting from applying the control sequence $\mathbf{u}_{0:N-1}$ to the system with dynamics $f$, initialized at the state $x_{0}$, is denoted by $x_{N}=f(x_{0},\mathbf{u}_{0:N-1})$, where each intermediate state is determined recursively with (1). We also denote a closed-loop trajecotry by $\mathbf{x}_{0:N}=\{x_{0},\dots,x_{N}\}$. This implies that there exists a feasible control sequence $\mathbf{u}_{0:N-1}$ such that $x_{k+1}=f(x_{k},u_{k})$ for all $k=0,\dots,N-1$.
@@ II-B Constrained Optimal Control Task
s.t., $x_{k+1}=f(x_{k},u_{k}),$
$y_{k}=h(x_{k})+n_{k},$
$x_{k}\in\mathcal{X},~{}u_{k}\in\mathcal{U},$
$\forall k\geq 0.$
@@
A closed-loop trajectory $\mathbf{x}_{0:N}=\{x_{0},\dots,x_{N}\}$ is called successful if $x_{N}\in\mathcal{X}_{f}$, and failed if $x_{N}\notin\mathcal{X}_{f}$.

### v2

@@ Abstract
Guaranteeing constraint satisfaction is challenging in imitation learning (IL), particularly in tasks that require operating near a system’s handling limits. Traditional IL methods, such as Behavior Cloning (BC), often struggle to enforce constraints, leading to suboptimal performance in high-precision tasks. In this paper, we present a simple approach to incorporating safety into the IL objective. Through simulations, we empirically validate our approach on an autonomous racing task with both full-state and image feedback, demonstrating improved constraint satisfaction and greater consistency in task performance compared to BC.
@@ II-A System
As a shorthand notation, we denote $\mathbf{u}_{0:N-1}=\{u_{0},\dots,u_{N-1}\}$ as a control sequence. The state at time NN resulting from applying the control sequence $\mathbf{u}_{0:N-1}$ to the system with dynamics $f$, initialized at the state $x_{0}$, is denoted by $x_{N}=f(x_{0},\mathbf{u}_{0:N-1})$, where each intermediate state is determined recursively with (1). We also denote a closed-loop trajecotry by $\mathbf{x}_{0:N}=\{x_{0},\dots,x_{N}\}$. This implies that there exists a feasible control sequence $\mathbf{u}_{0:N-1}$ such that $x_{k+1}=f(x_{k},u_{k})$ for all $k=0,\dots,N-1$.
While a formal observability analysis is not included in this work, we assume the system is observable.
@@ II-B Constrained Optimal Control Task
s.t., $x_{k+1}=f(x_{k},u_{k}),~y_{k}=h(x_{k})+n_{k},$
$x_{k}\in\mathcal{X},~u_{k}\in\mathcal{U},~~\forall k\geq 0,$
@@
A closed-loop trajectory $\mathbf{x}_{0:N}$ is called successful if $x_{N}\in\mathcal{X}_{f}$, and failed if $x_{N}\notin\mathcal{X}_{f}$.